OpenStudy (anonymous):

sue had 4 slices,tammy had 2 slices; 1/3 left over....how many slices did pizza have

OpenStudy (texaschic101):

sue had 4 slices tammy had 2 slices added, that is 6 slices. there was 1/3 left....that means 2/3 of pizza was already eaten. That means 6 slices is 2/3. 6/(2/3) = x/(3/3) -- 6 slices to 2/3 = x slices = 3/3 cross multiply 2/3x = (6)(3/3) 2/3x = 6 -- divide by 2/3 on both sides x = 6/(2/3) x = 6 * 3/2 x = 18/2 x = 9 check.. 2/3 of 9 = 2/3(9) = 18/3 = 6 slices 1/3 of 9 = 1/3(9) = 9/3 = 3 slices There were 9 slices of pizza in total.

OpenStudy (texaschic101):

do you have any questions ?